# Microquasar jet-cocoon systems as PeVatrons

### v1

@@ V Summary and Discussion
In this work, we studied the reacceleration and escape of CRs in microquasars within the framework of the shear acceleration mechanism. Our results showed that the CRs from microquasars could explain the observed proton energy spectrum above the PeV energy range. The predicted all-particle spectrum and composition are consistent with the LHAASO observations, where the relative fraction of difference nuclei is constrained by direct measurements around rigidity from 10 GV to 1 TV. The maximum acceleration energy of CRs from shear acceleration in microquasars is close to the knee around 3 PeV, and the escaping CRs have a hard spectrum.
@@
B.T.Z. is supported in China by National Key R&D program of China under the grant 2024YFA1611402. This work is partly supported by KAKENHI Nos. 22K14028, 21H04487, 23H04899 (SSK), and 20H05852 (KM). SSK acknowledges the support by the Tohoku Initiative for Fostering Global Researchers for Interdisciplinary Sciences (TI-FRIS) of MEXT’s Strategic Professional Development Program for Young Researchers. Numerical computations were in part carried out on Cray XD2000 at the Center for Computational Astrophysics, National Astronomical Observatory of Japan. K.M. is supported by the NSF Grant Nos. AST-2108466, AST-2108467 and AST-2308021. KM also acknowledges the SUGAR 2024 workshop in Madison, Wisconsin, which led to the idea of shear acceleration.
@@ Appendix A Dependence on source parameters
Here, we present the results for parameters that are different from the default one in Table 1, as indicated in Table 3. The predicted energy spectrum of escaping CRs is summarized in Fig. 5. The predicted CR intensity on Earth is presented in Fig. 6 and Fig. 7, respectively. We can see the peak energy predicted with Model-B is slightly below PeV, which is unable to explain the observed proton knee structure.
@@
Table 3: Model parameters used in Monte Carlo simulations.
@@
Figure 6: Similar to Fig. 3, but with parameters from Model-B.
Figure 7: Similar to Fig. 3, but with parameters from Model-C.
@@ Appendix B Propagate CRs in the Galaxy
In Fig. 8, we show the observed energy spectrum of CR species calculated from DRAGON code. The injection spectrum from Galactic SNRs is assumed to have broken powerlaw formula,
@@
where $\mathcal{R}_{\rm max}=0.4\rm~{}PV$. The adopted parameters are listed in  4 and the energy spectrum is presented in Fig. 8.
@@ H He C N O Ne
Table 4: Parameters used in DRAGON for the propagation of CRs from Galactic SNRs. Note the parameters for nuclei heavier than Neon is keep the same, where we not listed here.

### v2

@@ V Summary and Discussion
In this work, we studied the reacceleration and escape of CRs in microquasars within the framework of the shear acceleration mechanism. Our results showed that the CRs from microquasars could explain the observed proton energy spectrum above the PeV energy range. In our model, the relative fraction of nuclei is determined by the direct measurements of CRs. The predicted all-particle spectrum and composition ratio are consistent with the LHAASO observations if the CR injection rigidity is 100 GV to 1 TV. The maximum acceleration energy of CRs from shear acceleration in microquasars is close to the knee around 3 PeV, and the escaping CRs have a hard spectrum.
The required number of microquasars in our Galaxy is three to ten (compare Eqs. [5] and [7]). The lifetime of microquasars in our Galaxy is uncertain, but it might be 0.0310.03-1 Myr Yamamoto et al. (2008); Sakemi et al. (2023), which is comparable to or longer than the escape time of the Knee CRs ($\sim$80 kyr). If the lifetime is comparable or shorter than the CR escape time, the CR density in the ISM might be spatially inhomogeneous and temporary intermittent Giacinti and Semikoz (2023); Stall et al. (2025). In addition, we assume that microquasars in our Galaxy have identical CR spectral shape and CR luminosity for simplicity. In reality, there are distributions of CR luminosity and spectral shape, depending on the jet power and jet size of individual microquasars. Investigating the effect of these remains as a future work.
@@
We gratefully acknowledge the valuable discussions and insightful comments provided by Walter Winter, Martin Pohl, and Andrew Taylor. B.T.Z. is supported in China by National Key R&D program of China under the grant 2024YFA1611402. This work is partly supported by KAKENHI Nos. 22K14028, 21H04487, 23H04899 (S.S.K.), and 20H05852 (K.M.). S.S.K. acknowledges the support by the Tohoku Initiative for Fostering Global Researchers for Interdisciplinary Sciences (TI-FRIS) of MEXT’s Strategic Professional Development Program for Young Researchers. The Monte Carlo simulations of shear acceleration performed by S.S.K. were carried out on Cray XD2000 at the Center for Computational Astrophysics, National Astronomical Observatory of Japan. K.M. is supported by the NSF Grant Nos. AST-2108466, AST-2108467 and AST-2308021. K.M. also acknowledges the SUGAR 2024 workshop in Madison, Wisconsin, which led to the idea of shear acceleration.
@@ Appendix A Dependence on source parameters
Here, we present the results for parameters that are different from the default one in Table 1, as indicated in Table 3. The predicted energy spectrum of escaping CRs is summarized in Fig. 5. The predicted CR intensity on Earth is presented in Figs. 6 and Fig. 7, respectively. We can see the peak energy predicted with Model B is slightly below PeV, which is unable to explain the observed proton knee structure.
Table 3: Model parameters used in Monte Carlo simulations.
@@
Figure 6: Similar to Fig. 3, but with parameters from Model B.
Figure 7: Similar to Fig. 3, but with parameters from Model C.
@@ Appendix B Propagate CRs in the Galaxy
In Fig. 8, we show the observed energy spectrum of CR species calculated from DRAGON code. The injection spectrum from Galactic SNRs is assumed to have a broken powerlaw formula,
@@
where $\mathcal{R}_{\rm max}=0.4\rm~PV$. The adopted parameters are listed in Table 4 and the energy spectrum is presented in Fig. 8.
Table 4: Parameters used in DRAGON for the propagation of CRs from Galactic SNRs. Note that the parameters for the nuclei that were heavier than neon were keep the same, and not listed here.